Find a set of solutions for:
2(x-y)dx+(3x-y-1)dy=0

(2x+y-4)dx+(x-3y+12)dy=0

dy/dx=tanycotx - secycosx
1
Expert's answer
2020-04-06T07:58:23-0400

1) Find a set of solutions for

"2(x-y)dx+(3x-y-1)dy=0\\"

"y'={2(y-x) \above{2pt} 3x-y-1}"

Solving the system

"y_ \n0\n\u200b\t\n \u2212x_ \n0\n\u200b\t\n =0"

"3x_0-y_0-1=0"

"x_0={1 \above{2pt} 2}, y_0={1 \above{2pt} 2}"

We shall replace

"x=t+{1 \above{2pt} 2}"

"y=z+{1\above{2pt} 2}"

"y'=z'"

where t is a new independent variable,  z is a new function.Substitute into equation

"z'={2(z+ {1 \above{2pt} 2}-t- {1\above{2pt} 2}) \above{2pt} 3t+ {3 \above{2pt} 2} -z- {1 \above{2pt} 2}-1 }"

"z'={2(z-t) \above{2pt} 3t-z}"

we shall replace

"{z \above{2pt} t}=u(t),z=t.u. z'=tu'+u"

then,

"tu'+u={2(tu-t) \above{2pt} 3t-tu}"

"tu'+u={2u-2\above{2pt} 3-u}"

"tu'={2u-2 \above{2pt} 3-u}-u"

"tu'={2u-2-3u+u^2 \above{2pt} 3-u}"

"{3-u \above{2pt} (u+1)(u-2)} du={dt \above{2pt} t}"

"\int (-{4 \above{2pt} 3}{1 \above{2pt} u+1}+{1 \above{2pt} 3}{1 \above{2pt} u-2})du=\int {dt \above{2pt} t}"

"-{4 \above{2pt} 3}ln|u+1|+{1 \above{2pt} 3}ln|u-2|=ln|t|+{1 \above{2pt} 3} ln|c|"

"{(u-2) \above{2pt} (u+1)^4}=ct^3 ,u+1=0"

where "u={y- {1\above{2pt} 2}\above{2pt} x-{1 \above{2pt} 2} }"

The solution of the equation are

"{ {y- {1 \above{2pt} 2}\above{2pt} x- {1 \above{2pt} 2} } -2 \above{2pt}( {y- {1 \above{2pt} 2}\above{2pt} x- {1 \above{2pt} 2} } )^{4}+1 }" "=c(x-{1 \above{2pt} 2})^{3}"

"{y- {1\above{2pt} 2}\above{2pt} x-{1 \above{2pt} 2} }+1=0"

2) Find a set of solutions for

"(2x+y-4)dx+(x-3y+12)dy=0\\ y'=-\frac{2x+y-4}{x-3y+12}"

Solving the system

"2x_0+y_0-4=0"

"x\n_0-3y_0+12=0"

"x_0=0, y_0=4"

we shall replace

"x=t , y=z+4"

"y'=z'"

"z'=-{2t+z+4-4 \above{2pt} t-3z-12+12}"

"z'=-{2t +z\above{2pt} t-3z}"

we shall replace

"{z \above{2pt} t}=u(t)" ,"z=t.u, z'=tu'+u"

then

"tu'+u={2t+tu \above{2pt} 3tu-t}"

"tu'={2+u-3u^2+u \above{2pt} 3u-1}"

"-{3u-1 \above{2pt} 3u^2-2u-u} du={dt\above{2pt} t}"

"-{1 \above{2pt} 2}\int {d(3u^2-2u-2)\above{2pt} 3u^2-2u-2}du=\int{dt\above{2pt} t}"

"-{1\above{2pt} 2}ln|3u^2-2u-2|=ln|t|+{1 \above{2pt} 2}ln|c|"

"{1 \above{2pt} 3u^2-2u-2}=ct^2"

"3u^2-2u-2=0"

where "u={y-4\above{2pt} x}"

Then solution of the above equation are

"{1 \above{2pt} 3( {y-4 \above{2pt} x} )^{2}-2* {y-4\above{2pt} x}-2 }=cx^2"

"3({y-4 \above{2pt} x})^2-2*{y-4\above{2pt} x}-2=0"

3) Find a set of solutions for

"\frac{dy}{dx}=\tan y\cot x-\sec y\cos x\\ \frac{dy}{dx}=\frac{\sin y}{\cos y}\frac{\cos x}{\sin x}-\frac{1}{\cos y}\cos x\\ \frac{dy}{dx}=\frac{\sin y\cos x-\sin x \cos x}{\sin x\cos y}\\ (\sin x-\sin y)\cos x dx+\sin x \cos y dy=0"

"put" "u=\sin y, du=\cos ydy"

Substituting, we get

"(sinx\u2212u)cosxdx+sinxdu=0"

"(\sin x-u)\cos xdx+\sin x du=0\\ \frac{du}{dx}-u\cdot\frac{\cos x}{\sin x}=-\cos x"

The equation is first - order linear in u The integrating factor is

"I=e^{-\\int\\frac{\\cos x}{\\sin x}dx}=e^{-\\ln\\sin x}=\\frac{1}{\\sin x}\n\u200b"

Hence

"u\frac{1}{\sin x}=-\int\frac{\cos x}{\sin x}dx=-ln|\sin x|+c"

solve for u

"u=-\sin x\ln|\sin x|+c\sin x"

Put y back

"\sin y=-\sin x\ln|\sin x|+c\sin x"